For 12-13.

Big boss Egnaro decides to give each factory employee $1,000 each instead of fixing the slippery floors. The lives of these employees are always at risk with the odds of dying at 1 in 400. He has 10 employees. How much is the value of 1 statistical life in this situation?

$400,000

$4 million

$24,000

$60,000

$1,500

13.)

Use all the info in number 12 except now the number of employees = 80. How much is the value of 1 statistical life in this situation?

$400,000

$4 million

$10,000

$40,000

none of the above
Answer #1

12. Option 1. $400,000

Explanation:

The probability of death of 1 employe = 1/400 = 0.0025

Therefore, the value of 0.0025 statistical life is $1,000. So, the value of 1 statistical life is = $1,000/0/0025 = $400,000

13. Option 1. $400,000

Explanation: The value of 1 statistical life would remain fixed at $400,000 as the probability of 1 employe facing death remains the same at 0.0025 and the compensation paid to each employee remains the same as $1,000.